### v3

@@ 2.1 Hybrid Entropy Computing System
Coherent states quantum state of light closely resemble classical oscillation of light field, but also include quantum noise, specifically shot noise. In this hybrid system, single photon count of each time-bin experience Poisson noise which is taken advantage as a source of “entropy” or “fluctuation”, in each feedback loop. We define this parameter as “quantum fluctuation coefficient” that equals to $\frac{1}{\sqrt{N}}$, following the standard deviation of Poisson distribution where NN the number of photon collected in each time-bin. Therefore, throughout the measurement and feedback process, the injection of fluctuations enables a search of the solution space, which assists in bypassing the trapping of the system into local minima. The results section of this paper evidence solution quality of Dirac-3 affects by varying quantum fluctuation or shot noise. We expect that an all optical system could perform even better and we discuss a route to one in supplementary note 1. It requires many incremental loops to find the global minimum solutions to optimization problems with many variables. It is worth noting that since the optical parts of the system do not involve large scale non-Gaussian quantum superpositions, they could in principle be instead classically simulated, as has been argued in the coherent Ising machine setting [34, 35]. In spite of this fact, these machines still provide a useful demonstration of what can be done in optical computing, and a preview of what a fully optical system may be able to do. The ultimate goal of a quantum version of an entropy computer would be to take full advantage of the non-Gaussian statistics induced by Fock-state measurement, this is discussed in supplementary note 4.
@@ 3 Results
We first consider a simple two-variable optimization problem, defined through the objective function $g(x,y)=(x^{4}+y^{4})/4-5(x^{3}+y^{3})/3+3(x^{2}+y^{2})$. As depicted in Fig. 2(a), this function contains three local minima at $(\textup{x},\textup{y})=\{(0,3),(3,0),(3,3)\}$ as well as a global minimum at $(\textup{x},\textup{y})=(0,0)$. To solve this problem using Dirac-3, we consider 3 variables $v_{1}$, $v_{2}$, and $v_{3}$, with $v_{1}$ and $v_{2}$ representing $x$ and $y$ respectively, while $v_{3}$ representing a slack variable which is not coupled to any other variable but used to effectively relax the sum constraint. In this manner, one can consider a relatively large sum constraint of $R=100$ to incorporate a larger range of potential solutions. Figure 2(b) shows the evolution of the energy level for 2020 different runs as the system evolves toward an equilibrium point. The evolution of the variables versus the number of iterations are shown in Figs. 2(c) for three exemplary cases. These figures indicate rapid convergence of the solution after only a few iterations. Even more interestingly, the trajectories of the optimization variables in the two-dimensional $(x,y)$ plane (Figs. 2(d-k)) show that the system takes steps to avoid the local minima, even though in all cases the initial point was located within the attractor basins of the local minima.
@@
Now, considering a graph with NN nodes, described with adjacency matrix AA, i.e., $A_{ij}=1$ for adjacent nodes and $A_{ij}=0$ otherwise, the max-k-cut problem can be formulated as minimizing the following objective function: $E=\sum_{i,j}A_{ij}\vec{\mathbf{s}}_{i}\cdot\vec{\mathbf{s}}_{j}+\lambda\sum_{i}{\left\|\vec{\mathbf{s}}_{i}-\mathbf{1}\right\|}_{1}^{2}$, where the second term is an L1-norm regularization term with a parameter $\lambda$. It is straightforward to cast this objective function in the form of relation 1. This results in a quadratic function involving qNqN variable, where $q$ is the dimension of a single artificial spin state and NN is the size of the graph.
To test this formulation, we consider a randomly generated graph of $N=30$ nodes such that each two nodes are connected with a probability of $p=0.5$. Figure 4(a-c) visualizes exemplary partitioning of this graph into 22, 33 and 44 groups through the proposed entropy computing system. This particular graph has $233$ total links, and we found the maximum cut size to be $146$. Histograms of the cut sizes obtained for 100 different runs of the max-cut, max-3-cut, and max-4-cut problems using the entropy computing solver at schedule 1 are shown in Figs. 4(d-f), and using schedule 4 in Figs. 4(g-i). The results are compared with those obtained from Semi-Definite Programming (SDP) shown in Figs. 4(j-l). Here, we used the CVXPY package with the SCS solver to solve the semidefinite program. As expected, we observe that schedule 4 provides better results. In the version of Dirac-3 reported here, schedule 1 was set to have smaller number of feedback loops, lower quantum fluctuation coefficient, and higher mean photon number compare to schedule 4. In addition, as this figure clearly indicates, in all cases the entropy computing machine provides excellent results that outperform SDP results. As a reference, it is worth mentioning that SDP relaxation for max-cut has a performance guarantee of the ratio of $0.878$ [49], which translates to a cut size of $128$ for the particular graph considered here. For the results presented in Fig. 4, the regularization parameter was chosen to be $\lambda=5$.
@@ Acknowledgements
Concretely, we can map optimization problems to the interferometer mesh if relative phase is controlled, than interference between the modes can be used to encode an optimization problem, such that antiferromagnetic coupling is mapped to destructive interference, and ferromagnetic coupling to constructive. This is conceptually similar to a coherent Ising machine encoding but uses the amplitude degrees of freedom rather than phase degrees of freedom to store information. Since amplitude directly corresponds with photon number, Fock state measurements are natural in this setting. A natural way of storing a binary variable for example is in the relative amplitude/number of photons in two modes, such that more photons in one corresponds to 11 and in the other corresponds to 0, this is somewhat analogous to the encoding we currently use in the measurement-and-feedback setting.
@@
To map discrete optimization problems in the objective function of the form described in Eq. (4) of the main text, we allocate $k$ time bins to a discrete $k$-level variable. For example, in the case of binary variables, e.g., $s_{i}\in\{-1,1\}$, we consider two continuous variables to represent a binary variable as $\mathbf{s}_{i}=(x_{i},y_{i})^{t}$, which is a superpositon of the two states $(0,1)^{t}$ and $(1,0)^{t}$ with non-negative weights $x_{i}\geq 0$ and $y_{i}\geq 0$. Here, we consider penalty terms for penalizing results other than the discrete states above. In that case, as an example, the Max-Cut problem on a graph with adjacency matrix AA, can be formulated as finding state that minimizes the objective function $\sum_{i,j}A_{ij}s_{i}s_{j}$, which can be encoded in the following form:

### v4

@@ 2.1 Hybrid Entropy Computing System
Coherent states quantum state of light closely resemble classical oscillation of light field, but also include quantum noise, specifically shot noise. In this hybrid system, single photon count of each time-bin experience Poisson noise which is taken advantage as a source of “entropy” or “fluctuation”, in each feedback loop. We define this parameter as “quantum fluctuation coefficient” that equals to $\frac{1}{\sqrt{N}}$, following the standard deviation of Poisson distribution where $N$ the number of photon collected in each time-bin. Therefore, throughout the measurement and feedback process, the injection of fluctuations enables a search of the solution space, which assists in bypassing the trapping of the system into local minima. The results section of this paper evidence solution quality of Dirac-3 affects by varying quantum fluctuation or shot noise. We expect that an all optical system could perform even better and we discuss a route to one in supplementary note 1. It requires many incremental loops to find the global minimum solutions to optimization problems with many variables. It is worth noting that since the optical parts of the system do not involve large scale non-Gaussian quantum superpositions, they could in principle be instead classically simulated, as has been argued in the coherent Ising machine setting [34, 35]. In spite of this fact, these machines still provide a useful demonstration of what can be done in optical computing, and a preview of what a fully optical system may be able to do. The ultimate goal of a quantum version of an entropy computer would be to take full advantage of the non-Gaussian statistics induced by Fock-state measurement, this is discussed in supplementary note 4.
@@ 3 Results
We first consider a simple two-variable optimization problem, defined through the objective function $g(x,y)=(x^{4}+y^{4})/4-5(x^{3}+y^{3})/3+3(x^{2}+y^{2})$. As depicted in Fig. 2(a), this function contains three local minima at $(\textup{x},\textup{y})=\{(0,3),(3,0),(3,3)\}$ as well as a global minimum at $(\textup{x},\textup{y})=(0,0)$. To solve this problem using Dirac-3, we consider 3 variables $v_{1}$, $v_{2}$, and $v_{3}$, with $v_{1}$ and $v_{2}$ representing $x$ and $y$ respectively, while $v_{3}$ representing a slack variable which is not coupled to any other variable but used to effectively relax the sum constraint. In this manner, one can consider a relatively large sum constraint of $R=100$ to incorporate a larger range of potential solutions. Figure 2(b) shows the evolution of the energy level for $20$ different runs as the system evolves toward an equilibrium point. The evolution of the variables versus the number of iterations are shown in Figs. 2(c) for three exemplary cases. These figures indicate rapid convergence of the solution after only a few iterations. Even more interestingly, the trajectories of the optimization variables in the two-dimensional $(x,y)$ plane (Figs. 2(d-k)) show that the system takes steps to avoid the local minima, even though in all cases the initial point was located within the attractor basins of the local minima.
@@
Now, considering a graph with $N$ nodes, described with adjacency matrix $A$, i.e., $A_{ij}=1$ for adjacent nodes and $A_{ij}=0$ otherwise, the max-k-cut problem can be formulated as minimizing the following objective function: $E=\sum_{i,j}A_{ij}\vec{\mathbf{s}}_{i}\cdot\vec{\mathbf{s}}_{j}+\lambda\sum_{i}{\left\|\vec{\mathbf{s}}_{i}-\mathbf{1}\right\|}_{1}^{2}$, where the second term is an L1-norm regularization term with a parameter $\lambda$. It is straightforward to cast this objective function in the form of relation 1. This results in a quadratic function involving $qN$ variable, where $q$ is the dimension of a single artificial spin state and $N$ is the size of the graph.
To test this formulation, we consider a randomly generated graph of $N=30$ nodes such that each two nodes are connected with a probability of $p=0.5$. Figure 4(a-c) visualizes exemplary partitioning of this graph into $2$, 33 and $4$ groups through the proposed entropy computing system. This particular graph has $233$ total links, and we found the maximum cut size to be $146$. Histograms of the cut sizes obtained for 100 different runs of the max-cut, max-3-cut, and max-4-cut problems using the entropy computing solver at schedule 1 are shown in Figs. 4(d-f), and using schedule 4 in Figs. 4(g-i). The results are compared with those obtained from Semi-Definite Programming (SDP) shown in Figs. 4(j-l). Here, we used the CVXPY package with the SCS solver to solve the semidefinite program. As expected, we observe that schedule 4 provides better results. In the version of Dirac-3 reported here, schedule 1 was set to have smaller number of feedback loops, lower quantum fluctuation coefficient, and higher mean photon number compare to schedule 4. In addition, as this figure clearly indicates, in all cases the entropy computing machine provides excellent results that outperform SDP results. As a reference, it is worth mentioning that SDP relaxation for max-cut has a performance guarantee of the ratio of $0.878$ [49], which translates to a cut size of $128$ for the particular graph considered here. For the results presented in Fig. 4, the regularization parameter was chosen to be $\lambda=5$.
@@ Acknowledgements
Concretely, we can map optimization problems to the interferometer mesh if relative phase is controlled, than interference between the modes can be used to encode an optimization problem, such that antiferromagnetic coupling is mapped to destructive interference, and ferromagnetic coupling to constructive. This is conceptually similar to a coherent Ising machine encoding but uses the amplitude degrees of freedom rather than phase degrees of freedom to store information. Since amplitude directly corresponds with photon number, Fock state measurements are natural in this setting. A natural way of storing a binary variable for example is in the relative amplitude/number of photons in two modes, such that more photons in one corresponds to $1$ and in the other corresponds to 0, this is somewhat analogous to the encoding we currently use in the measurement-and-feedback setting.
@@
To map discrete optimization problems in the objective function of the form described in Eq. (4) of the main text, we allocate $k$ time bins to a discrete $k$-level variable. For example, in the case of binary variables, e.g., $s_{i}\in\{-1,1\}$, we consider two continuous variables to represent a binary variable as $\mathbf{s}_{i}=(x_{i},y_{i})^{t}$, which is a superpositon of the two states $(0,1)^{t}$ and $(1,0)^{t}$ with non-negative weights $x_{i}\geq 0$ and $y_{i}\geq 0$. Here, we consider penalty terms for penalizing results other than the discrete states above. In that case, as an example, the Max-Cut problem on a graph with adjacency matrix $A$, can be formulated as finding state that minimizes the objective function $\sum_{i,j}A_{ij}s_{i}s_{j}$, which can be encoded in the following form: